OpenStudy (emilyjones284):

im very confused on this one..ill give a medal (:

OpenStudy (anonymous):

MEDAL YUMYUM

OpenStudy (anonymous):

LOL :D

OpenStudy (jdoe0001):

answer is 2 bushels of corn!!

OpenStudy (emilyjones284):

\[(4x+10)^\frac{ 1 }{ 4 }=(9+7x)^\frac{ 1 }{ 4 }\]

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

raise both sides by "4 power" and all you're left with is 4x+10 = 9+7x

OpenStudy (campbell_st):

raise each expression to a power of 4 it will eliminate the fractional indices the you are solving 4x + 10 = 9 + 7x subtract 4x from both sides of the equation then subtract 9 from both sides of the equation... then you should be able to find the answer

OpenStudy (emilyjones284):

oh ok i got it thanks (: